What is the value of this expression when x =4 and y= -3
x^3+y^2+x^2y^3- (x+y)

Answer by algebrahouse.com(1659)

x³ + y² + x²y³ - (x + y), when x = 4 and y = -3
= 4³ + (-3)² + (4)²(-3)³ - [4 + (-3)] {substituted 4 for x and -3 for y}
= 64 + 9 + 16(-27) - (4 - 3) {simplified}
= 64 + 9 - 432 - 1 {multiplied}
= -360 {added and subtracted}

For more help from me, visit: www.algebrahouse.com